OpenStudy (anonymous):

Simplify the expression completely. (assume all variables are positive) sqrt (9a^12 b^-10)

OpenStudy (psymon):

Basically, if an exponent is negative, flip it upside down. For example: \[x ^{-10}=\frac{ 1 }{ x ^{10} }\] \[\frac{ 3 }{ x ^{-5} }=3x ^{5} \]Basically you can do that with your negative exponent. You just move it to the bottom to make it positive.

OpenStudy (anonymous):

I tried this as an answer 3a sqrt 1/b^10 and still did not accept that answer. Am I doing something wrong

OpenStudy (psymon):

I apologize, I didn't see that you wrote sqrt x_x

OpenStudy (psymon):

\[\sqrt{9a ^{12}b ^{-10}}\]correct?

OpenStudy (anonymous):

\[3a \sqrt{b^-10}\] is the answer I got but wont accept it I assume because the whole (assume all vairiables are positive). Yes the forumula you pu is correct

OpenStudy (psymon):

Alright, awesome. Well the first thing to know is what I wrote above, about how you can move negative exponents around. So I can start by rewriting the problem like this: \[\sqrt{\frac{ 9a ^{12} }{ b ^{10} }}\]Kinda understand this part?

OpenStudy (anonymous):

yes, this way there are no negative exponents

OpenStudy (psymon):

Alright, cool. Now this is just to make sure, but do we know: \[\sqrt{x ^{2}}\] and what that is? It may sound lke a stupid question, but I just want to be sure.

OpenStudy (anonymous):

\[x ^{1/2}\]?

OpenStudy (psymon):

It would just be x. \[\sqrt{x}=x ^{\frac{ 1 }{ 2 }}\], but since we have that x^2 under the root, it just reduces to x.

OpenStudy (anonymous):

ok makes sense

OpenStudy (psymon):

Alright. So the next bit of important info is this: \[\sqrt{x ^{2}}=\sqrt{x}*\sqrt{x}\]If you know how to factor, we can actually do this. Another example: \[\sqrt{18}=\sqrt{2}*\sqrt{3}*\sqrt{3}\] That make sense?

OpenStudy (anonymous):

yes

OpenStudy (psymon):

Okay, cool. So if we can factor something inside of the square root, we can break it apart like this. So let's do that with your problem. I'llstart with the top and rewrite it: \[\sqrt{9a ^{12}}=\sqrt{9}*\sqrt{a ^{2}}*\sqrt{a ^{2}}*\sqrt{a ^{2}}*\sqrt{a ^{2}}*\sqrt{a ^{2}}*\sqrt{a ^{2}}\]You see why?

OpenStudy (anonymous):

yes 9 is sqrt alone then we just break down the exponent of a^12

OpenStudy (psymon):

Correct. So all 7 of those square roots can simply. We get 9 to become 3 and those other 6 all become a. So that means it'll simplify to: \[3a ^{6} \]

OpenStudy (anonymous):

and bottom would break down to \[\sqrt{b10}*\sqrt{b10}*\sqrt{b10}*\sqrt{b10}*\sqrt{b10}\]

OpenStudy (psymon):

This can also be done with the b on the bottom. \[\sqrt{b ^{10}}=\sqrt{b ^{2}}*\sqrt{b ^{2}}*\sqrt{b ^{2}}*\sqrt{b ^{2}}*\sqrt{b ^{2}} \]

OpenStudy (anonymous):

ment 2 right lol copy past messed up so \[\sqrt{b^5}\]

OpenStudy (psymon):

Well, b^5, but no root. Each of those 5 b^2 roots just become b, so you have b^5 on bottom to finally get: \[\frac{ 3a ^{6} }{ b ^{5} }\]

OpenStudy (psymon):

If you notice, when the power under a square root is even, you can just divide it by 2 and that will be how many you have outside of the root. So even something like: \[\sqrt{x ^{1204}}=x ^{602}\]can easily be done, just divide by 2 when you see the power is even.

OpenStudy (anonymous):

Oh ok I got it yeah not flipping it inside the sqrt is the main thing I was doing, thank you so much

OpenStudy (psymon):

yeah, np :3